OpenStudy (anonymous):

Which statement is true about the product square root of 2(3square root of 2 + square root of 14)? It is rational and equal to 12. It is rational and equal to 18. It is irrational and equal to 6 + 2square root of 18. It is irrational and equal to 3 +square root of 18.

OpenStudy (anonymous):

Help plz need answer fast no time :(

OpenStudy (anonymous):

I give medal :(!

OpenStudy (anonymous):

Everyone who answer will get medal

OpenStudy (anonymous):

:D

OpenStudy (anonymous):



OpenStudy (solomonzelman):

\[2(3\sqrt{2}+\sqrt{14})\]this?

OpenStudy (anonymous):

+ 18 on end sorry

OpenStudy (solomonzelman):

\[2(3\sqrt{2}+\sqrt{14}+18)\]

OpenStudy (anonymous):

just 18

OpenStudy (anonymous):

no 14

OpenStudy (solomonzelman):

\[2(\sqrt{3}+\sqrt{18})\]

OpenStudy (anonymous):

2(3 squareroot 2 + 18

OpenStudy (anonymous):

that is real one

OpenStudy (solomonzelman):

\[3(\sqrt{2}+18)\]

OpenStudy (solomonzelman):

right?

OpenStudy (anonymous):

:Facepalm:

OpenStudy (solomonzelman):

\[2(\sqrt{3}+18)\]

OpenStudy (anonymous):

TWO(THREE square root2 + 18)

OpenStudy (anonymous):

thats real one k?

OpenStudy (solomonzelman):

\[2(3\sqrt{2}+18)\]

OpenStudy (anonymous):

Yes :D

OpenStudy (solomonzelman):

Nice!!!!!!

OpenStudy (anonymous):

Can you slove me answer?

OpenStudy (solomonzelman):

\[a(b+c)=ab+ac\]\[2(3\sqrt{2}+18)=(2 \times 3\sqrt{2})~~~+~~~(2 \times 18)=6\sqrt{2}+36\]

OpenStudy (anonymous):

ok then whats next

OpenStudy (solomonzelman):

that's it, I expanded it as far as possible.

OpenStudy (anonymous):

my answer choices say 6 squareroot 2 + 18

OpenStudy (anonymous):

you say 6 squareroot 2 + 36 but it no in my choices